If you cut out the largest possible circular pie crust for a square piece of dough that is "d" inches on a side, what is the area of the remaining dough? Express your answer as polynomial in "d".
Answer by mananth(12270)

area of square piece = d*d= d^2
...
area of circle = pi*r^2 where r= radius.
....
diameter = d
radius = d/2
area = pi(d/2 )^2
area = pi*d^2/4
..
remaining area = area of square - arae of circle
d^2- pi*d^2/4